OpenStudy (anonymous):

Hi, i really need help with this, i just can't seem to fully understand this, and i need this to help improve my grade. its really important. 22. Write the equation for the vertical line that contains point E(10, –3).

OpenStudy (lukebluefive):

Vertical lines are of the form: \[x = constant\] Point coordinates are usually arranged in the following way: \[(x, y)\] So the vertical line that contains the point (10, -3) must have the following equation: \[x = 10\]

OpenStudy (lukebluefive):

Here is an approximate graph of the situation: 

OpenStudy (anonymous):

Did you just give me the answer?

OpenStudy (anonymous):

thank you so much

OpenStudy (lukebluefive):

You're welcome, hope you understand it better now.

OpenStudy (anonymous):

i do, thanxs

OpenStudy (anonymous):

hey can you help me with something else

OpenStudy (lukebluefive):

Sure?

OpenStudy (anonymous):

21. Find the values of x, y, and z. The diagram is not to scale.

OpenStudy (anonymous):

nbbj

OpenStudy (lukebluefive):

Assuming the values on the diagram are degrees, the following can be said about x, based on the definition of a triangle:

OpenStudy (lukebluefive):

\[x = 180 - (63 + 36) = 180 - 99 = 81\]

OpenStudy (lukebluefive):

Wait... is 36 the angle for the entire top?

OpenStudy (anonymous):

yes

OpenStudy (lukebluefive):

What does the 13 on the side belong to?